Solve for the base. Round to hundredths when necessary.

What is the base if the rate is 16.9% and the portion is 454?


1
Expert's answer
2021-02-24T07:25:02-0500

"rate=portion/base \times 100\%"

"16.9\%=454/base \times 100\%"

"16.9\/100=454\/base"

"base =(454\times 100)/16.9=2,686.39"